OpenStudy (anonymous):

What is the volume of a cube with a side of length 14 units? A. 2744 cubic units B. 196 cubic units C. 84 cubic units D. 42 cubic units

jigglypuff314 (jigglypuff314):

same idea as the previous one :) volume of cube = (side)^3 = (side)*(side)*(side)